if the length of each side of a square is increased by 4 inches, the area of the square becomes 9 times greater. Find the perimeter of the original square
Answer by lwsshak3(11628)

if the length of each side of a square is increased by 4 inches, the area of the square becomes 9 times greater. Find the perimeter of the original square.
**
let x=original length of each side of a square
x^2=area of original square
(x+4)=side increased by 4
(x+4)^2=area of larger square
(x+4)^2=9x^2
x^2+8x+16=9x^2
8x^2-8x-16=0
x^2-x-2=0
(x-2)(x+1)=0
x=-1(reject)
or
x=2
perimeter of the original square=4x=8 inches